OpenStudy (anonymous):

Give examples of two numbers a and b such that(a+b)^2=a^2+b^2 and explain why the equation is only true in those cases.

OpenStudy (ddcamp):

(a + b)² = a² + 2ab + b² So when does a² + 2ab + b² = a² + b²?

OpenStudy (anonymous):

Can you explain more.

OpenStudy (ddcamp):

a² + 2ab + b² = a² + b² -(a² + b²) -(a² + b²) 2ab = 0 ab = 0 a = 0 or b = 0